it takes one man 3 hours to pick up all the trash on a highway. it takes his partner 5 hours. if they work together, how long would it take?
Answer by Alan3354(21605)

it takes one man 3 hours to pick up all the trash on a highway. it takes his partner 5 hours. if they work together, how long would it take?
-----------------------
It's obvious it's less than 3 hours. And, if they both did it in 3 hours, it would be 1.5 hours, so it'll be between those 2 numbers.
You have to use the inverses, not 3 hours to do it, but how much is done per hour.
In 1 hour, the 1st man does 1/3 of the job, the 2nd does 1/5.
Together, they do 1/3 + 1/5 = 8/15 of the job.
So the whole job takes 15/8 hours, to 1.875 hours, which is between 1.5 and 3.
-------------
A quicker way to do these is product/sum: 3*5/(3+5) = 15/8